OpenStudy (anonymous):

I need to know the steps for doing this word problem.A number is 6 times another number. When the numbers are aded their sum is 49 what are the numbers

myininaya (myininaya):

so do you have x=6y and 6y+x=49

myininaya (myininaya):

oops its x+y=49

OpenStudy (radar):

x = one number 6x = the other number 6x+x=49 7x=49 x =7 other number is 6 times 7 or 42

myininaya (myininaya):

right!

OpenStudy (anonymous):

omg thanks